OpenStudy (anonymous):

If there is one chance in five of picking a red marble, then what is the probability of not picking a red marble? Write the probability as a fraction and decimal. Can someone please help me?

OpenStudy (anonymous):

1-(1/5) = 4/5=0.8

OpenStudy (texaschic101):

I agree...if there is 1 chance in 5 of picking a red marble, then there is 4 chances in 5 of NOT picking a red marble. 4 chances out of 5 = 4/5 which equals 0.8